OpenStudy (anonymous):

Quadrilateral ABCD is similar to quadrilateral WXYZ. If AD = 12, DC = 6, and ZY = 36, find WZ.

OpenStudy (anonymous):

here WZ/AD = YZ/DC or WZ = 12 *36/6 =72

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome dear